OpenStudy (anonymous):

sqrt 3 times sqrt 21 equals?

OpenStudy (anonymous):

\[\sqrt{3} \times \sqrt{21} = sqrt{63}\]

OpenStudy (anonymous):

Sorry, \[\sqrt{3} \times \sqrt{21} = \sqrt{63}\]

OpenStudy (hoblos):

\[3\sqrt{7}\]

OpenStudy (anonymous):

\[\sqrt{3} \times \sqrt{21} = \sqrt{21 \times 3} = \sqrt{63}\]

OpenStudy (anonymous):

3squareroot7

OpenStudy (anonymous):

\[3\sqrt{7}\]

OpenStudy (anonymous):

\[\sqrt{3\sqrt{7?}}\]